7. Differentiate the function x = y ^ 4 - 2y ^ 3


8. Differentiate the function t = 1/2 * t ^ 4 - 5t - 3


9. Differentiate the function y = (x ^ 2 - 2) ^ 2


10. Differentiate the function y = 1/(x + 7)


11. Find the slope of function y = 1/(x ^ 2), 12, 1/4 12. Find the slope of function y^ 2 =4x.(1,2)


13. Find the slope of function y = 1/(x + 1), (- 2, 1)


1
Expert's answer
2022-03-22T18:37:04-0400

"1) x=y^4-2y^3\\
1=4y^3\cdot y'-6y^2\cdot y'\\
1=y'(4y^3-6y^2)\\
2) y=\frac{1}{2}t^4-5t-3\\
y'=\frac{1}{2}\cdot 4 t^3-5=2t^3-5\\
3) y=(x^2-2)^2\\
y'=2(x^2-2)\cdot 2x=4x(x^2-2)\\
4)y=\frac{1}{x+7}=(x+7)^{-1}\\
y'= - (x+7)^{-2}= \frac{-1}{(x+7)^2}\\
5) y=\frac{1}{x^2}=x^{-2}, (12,\frac{1}{4})\\
y'=k=\tan \alpha=f'(x_0)\\
y'=-2x^{-3},\\
y'(12)=-2(12)^{-3}=-\frac{1}{864}\\
\tan \alpha=-\frac{1}{864}\\
y'(\frac{1}{4})=-2(\frac{1}{4})^{-3}=-\frac{1}{32}\\
\tan \alpha=-\frac{1}{32}\\
6)y^2=4x, (1,2)\\
y=\sqrt{4x}\\
y'=\frac{1}{2\sqrt{4x}}\cdot 4=\frac{1}{\sqrt{x}}\\
y'(1)=\frac{1}{\sqrt{1}}=1\\
\tan\alpha=1\\
 \alpha=45^0\\
7) y=\frac{1}{x+1}, (-2,1)\\
y'=-\frac{1}{(x+1)^2}\\
y'(-2)=-\frac{1}{(-2+1)^2}=-1\\
\tan \alpha=-1\\
 \alpha=135^0\\"